Richard has just been given a 10-question multiple-choice quiz in his history class. Each question has four answers, of which only one is correct. Since Richard has not attended class recently, he doesn't know any of the answers. Assuming that Richard guesses on all ten questions, find the indicated probabilities. (Round your answers to three decimal places.)

What is the probability that he will answer all the questions correctly the answer is 0.000

My question is how did they get that answer
Answer #1

Solution: Given data 25, n = 10 p=4 q=1-1/4 = 3/4 that he correctly is distribution To determine the probability will answer

Please like if it helps me..

Thank you...